OpenStudy (anonymous):

Simplify. Express the quotient as a radical expression. eleventh root of x to the tenth power over eleventh root of x to the sixth power

OpenStudy (anonymous):

is this is your question? \[\Large \frac{\sqrt[11]{x^{10}}}{\sqrt[11]{x^6}}\]

OpenStudy (anonymous):

@gioanny39x

OpenStudy (anonymous):

yes :)

OpenStudy (anonymous):

@sami-21

OpenStudy (anonymous):

ok can be written under one radical as \[\Large \sqrt[11]{\frac{x^{10}}{x^6}}\] now if the basis are same power get subtracted or added using the fractions exponents rule \[\Large \frac{x^a}{x^b}=x^{a-b}\] if (b is less than a) \[\Large \sqrt[11]{x^{10-6}}\] or \[\Large \sqrt[11]{x^4}\] or \[\Large (x)^{\frac{4}{11}}\] i hope it is clear!